### v1

# Klear: Unified Multi-Task Audio-Video Joint Generation]Kling Team, Kuaishou Technology

Abstract
Audio–video joint generation has progressed rapidly, yet substantial challenges still remain. Non-commercial approaches still suffer audio-visual asynchrony, poor lip–speech alignment, and unimodal degradation, which can be stemmed from weak audio–visual correspondence modeling, limited generalization, and scarce high-quality dense-caption data. To address these issues, we introduce Klear and delve into three axes—model architecture, training strategy, and data curation. Architecturally, we adopt a single-tower design with unified DiT blocks and an Omni-Full Attention mechanism, achieving tight audio–visual alignment and strong scalability. Training-wise, we adopt a progressive multitask regime—random modality masking to joint optimization across tasks, and a multistage curriculum, yielding robust representations, strengthening A-V aligned world knowledge, and preventing unimodal collapse. For datasets, we present the first large-scale audio–video dataset with dense captions, and introduce a novel automated data-construction pipeline which annotates and filters millions of diverse, high-quality, strictly aligned audio–video–caption triplets. Building on this, Klear scales to large datasets, delivering high-fidelity, semantically and temporally aligned, instruction-following generation in both joint and unimodal settings while generalizing robustly to out-of-distribution scenarios. Across tasks, it substantially outperforms prior methods by a large margin and achieves performance comparable to Veo 3, offering a unified, scalable path toward next-generation audio–video synthesis.
1 Introduction
The Kinetoscope was the world’s first film projector, invented by Thomas Edison in 1891. The Kinetophone, on the other hand, was the first sound film projector, combining the Kinetoscope with synchronized sound playback technology, marking a major breakthrough in sound film technology. Accordingly, vision and auditory are complementary and jointly indispensable modalities for depicting the real world. Consequently, audio–video joint generation has emerged as one of the most prominent trends in generative AI. Commercial systems such as Veo 3 and Sora 2 have achieved impressive performance with strong semantic alignment, while diverse open-source models with varying architectures are also emerging. Nevertheless, research on joint audio–visual generation remains nascent, and current models—even some commercial systems—still exhibit audio–visual asynchrony, lip–speech mismatches, and degradation in unimodal quality. We attribute several factors: (1) Architecture. Most T2AV models like JavisDiT [1], UniVerse-1 [2] and Ovi [3] employ a single-tower architecture with a cross-attention module, resulting in limited audio-video interaction and alignment. (2) Training strategies and data. Currently, mainstream methods perform single-task training, which might bring about biased representations and struggle to exploit underlying audio-video correlation and world knowledge.
To this end, informed by the above observations, we identify several key contributing factors. (1) Architecturally, prevailing designs hinder thorough cross-modal interaction: most existing models adopt dual-tower architectures with modality-specific initialization, learn each modality independently, and rely on shallow fusion via cross-attention or adapters, which fails to fully align audio–visual features. (2) For data construction, there is a pronounced lack of diverse, high-quality, and densely annotated audio–video aligned generation datasets, as well as scalable, high-quality annotation methodologies for constructing them. (3) From a learning-strategy perspective, most existing methods are trained exclusively on text-to-AV generation. This single-task regime induces overfitting and representation bias, which in turn hinders generalization and degrades unimodal performance.
To address these issues, we propose Klear, which introduces coordinated improvements at the architectural, learning-strategy, and data levels. Specifically, on the architectural side, we adopt a single-tower backbone with unified DiT blocks. Each block integrates an Omni-Full Attention module that jointly attends to four streams—audio, audio captions, video, and video captions—thereby facilitating cross-modal fusion and interaction, achieving tight A-V alignment and stronger coupling to textual conditions, which offers a higher scaling ceiling. For training, we design a progressive multitask training regime: random modality masking sustains joint optimization over T2AV/TI2AV/TI2V/T2V/T2A. A performance-adaptive pretrain–post train curriculum tunes data mixtures and quality, yielding robust, generalizable representations that exploit A/V correlations and world knowledge. From the data construction, we introduce an automated annotation pipeline that enables efficient model scaling. Building on these three insights, our model achieves strong performance in both joint and unimodal generation, delivering high fidelity, tight audio–visual alignment, natural outputs, and robust instruction following with favorable scaling behavior. On the Verse-Bench, it surpasses prior methods by a large margin and generalizes well to out-of-distribution (OOD) scenarios. Our main contributions are summarized as follows:
  • We introduce Klear, a unified framework for multi-task audio–video joint generation that, to our knowledge, is the first model to achieve performance comparable to Veo 3. Klear effectively resolves semantic and temporal audio–visual misalignment while delivering high-fidelity generation.
  • Our key technical novelties lie in the unified single-tower architecture with the omni-full attention mechanism for seamless audio–visual fusion, and a progressive multi-task training strategy that promotes generalizable representations and prevents unimodal performance degradation.
  • We propose a large-scale, high-quality audio-video dataset consisting of 81 million samples with accurate dense captions, along with an automated data generation pipeline and a high-quality audio-video generation dataset.
  • Extensive experiments demonstrate that Klear consistently excels in both unimodal and joint audio–video generation, outperforming prior state-of-the-art on the unimodal benchmark and the AV joint-generation benchmark consistently.
2 Related Works
Text-to-Video Generation (T2V). Diffusion models have revolutionized video generation, with AnimateDiff [4] and Video Diffusion Models [5] leading the way. Stable Video Diffusion [6] emphasized the importance of large, high-quality datasets for performance. Early models used U-Net backbones, but Sora [7] introduced the Diffusion Transformer (DiT) architecture [8], trained on extensive video corpora. Open-source models like CogVideoX [9], HunyuanVideo [10], and the WAN series [11] joined closed-source systems like Kling [12] and Veo 2 [13]. Most models share a common architecture: a 3D VAE compresses videos into spatiotemporal latents, and a DiT performs denoising. Data quality and scale remain crucial for model success, highlighting the importance of data curation and processing.
Image-to-Video Generation (I2V). Early I2V systems extended T2V by conditioning on a single frame via latent concatenation or CLIP-based feature injection [14]. Later, cascaded designs like I2VGen-XL [15] and dual-injection frameworks like DynamiCrafter [16] introduced structured pipelines for better motion and appearance fidelity. Adapter-based approaches like LAMP [17] and I2V-Adapter [18] integrate cross-frame attention. Recent works on motion modeling and stable sampling include Motion-I2V [19] and FrameBridge [20]. Despite these advancements, I2V still faces challenges with curated datasets, limited long-range motion modeling, and the trade-off between appearance fidelity and motion realism.
Text-to-Audio Generation (TTA). Recent advances in generative models have propelled text-to-audio generation. Models like Make-An-Audio [21] and AudioLDM [22] synthesize audio via iterative denoising of text-conditioned latent representations. Tango [23], Audio Flamingo [24], and others expand latent spaces and enhance cross-modal alignment. Stable Audio [25] employs hierarchical latent diffusion for high-fidelity output. AudioStory [26] introduces a unified generation framework, achieving long-form audio generation for the first time, while flow-matching techniques like VoiceBox [27] enable zero-shot style transfer. TangoFlux [28] optimizes text-audio alignment with CLAP-ranked preferences. Though these methods excel at semantic alignment, they are limited to short durations and lack flexibility for complex, evolving instructions, underscoring the need for TTA models that handle long, complex tasks.
Audio-Video Joint Generation (T2AV). Pioneering efforts like MM-Diffusion [29] use coupled U-Net backbones, while DiT-based approaches dominate. AV-DiT [30] adapts pre-trained image DiTs with lightweight adapters, and UniForm [31] uses a unified single-tower architecture for audio-video tokens. Key challenges include precise spatio-temporal synchronization, addressed by methods like JavisDiT [1] (hierarchical prior), Ovi [3] (twin-backbone design), and SyncFlow [32] (dual-DiT with Rectified Flow Matching). Other research orchestrates unimodal experts, such as MMDisCo [33] and Universe-1 [2], which combine specialized models at the block level. Despite advances in architecture and data, most models focus on sound effects or music, leaving synchronized speech and video synthesis an underexplored challenge.
3 Klear
Figure 1: Overview of Klear. The model takes four inputs: video, video-related text, audio-related text, and audio. Each input is individually encoded by respective encoders, then fed into the MM-DiT. The MM-DiT module outputs the latent variables of video and audio, which are then decoded separately into video and audio.
3.1 Preliminary
Problem Definition. Our goal is to enable the generation of both audio and video within a single model, given various prior conditions. We denote the denoising network as $\epsilon_{\theta}(\cdot)$, the text condition as $c$. Let $\{z^{a}_{t}\}_{t\in[0,1]}$ and $\{z^{v}_{t}\}_{t\in[0,1]}$ denote the latent variables at timestamp $t$ for audio and video, respectively. Here, $t=0$ denotes the final timestamp of pure Gaussian noise. During inference, $\epsilon_{\theta}(\cdot)$ recursively performs denoising from $t=0$ to $t=1$ to produce the final generation, $\hat{z}^{a}_{1},\hat{z}^{v}_{1}$, as shown below:
$\hat{z}^{a}_{t^{\prime}},\hat{z}^{v}_{t^{\prime}}=\epsilon_{\theta}(z^{a}_{t},z^{v}_{t},t,c),$ (1)
Conditional Flow-Matching. We employ flow matching as the denoising objective. The model needs to learn the velocity field that transforms pure noise $p_{0}=\mathcal{N}(0,\boldsymbol{I})$ to the underlying data distribution $p_{\text{data}}$. In practice, we perform linear interpolation $x_{t}=(1-t)x_{0}+tx_{1}$ to construct a distribution at timestamp $t$. Here, $x_{0}\sim p_{0}$ and $x_{1}\sim p_{\text{data}}$. Given the condition $c$, the model $\epsilon_{\theta}(\cdot)$ is trained to predict the target velocity, i.e., constantly as $u_{t}=x_{1}-x_{0}$:
$\mathcal{L}_{\text{FM}}$ $=\mathbb{E}_{t,c,{x}_{0},{x}_{1}}\Big\|({x}_{1}-{x}_{0})-\epsilon_{\theta}\big(t{x}_{1}+(1-t){x}_{0},t,c\big)\Big\|_{2}^{2},$ (2)
$\text{where}\quad t\sim\mathcal{U}(0,1),\ {x}_{0}\sim\mathcal{N}(0,\boldsymbol{I}),\ {x}_{1}\sim p_{\text{data}}.$
Latent Encoding. The model takes four inputs: video, video-related text, audio-related text, and audio, where video-related text represents the video caption and audio-related text represents the audio caption and speech text. Video is encoded by the 3d casual visual encoder from CogVideoX [9], We use Qwen3-8B Embedding [34] as the encoder for audio and video captions.
3.2 Single Tower with Full Attention
Single Tower DiT. To ensure a thorough audio-video fusion, we employ a single-tower architecture. As shown in Fig. 1, following Stable Diffusion 3 [35], we employ Multimodal Diffusion (MMDiT) to take the sequences of all modalities as input and perform full attention. Specifically, there are four inputs, i.e., video, video-related text, audio-related text, and audio. Each type of input is individually encoded into latents with respective encoders, then fed into the MM-DiT. The MM-DiT module outputs the latent variables of video and audio in two streams, which are then decoded separately to perform video and audio generation.
Mixed Dimension Rotary Position Embedding (MixD-RoPE). Another key architectural innovation is Mixed Dimension Rotary Position Embedding (MixD-RoPE). As shown in Fig. 1 (d), to enhance the positional information introduced by various aspect ratios and duration in videos, we apply 3D RoPE encoding across three dimensions, i.e., temporal, width and height for video embedding. This 3D RoPE incorporates both absolute and relative position dependency in videos. For audio modality, we employ compatible temporal 1D positional encodings, while its position number is initialized by incrementing the maximum temporal position ID of the video modality by one. As a result, we build MixD-ROPE with a shared temporal position ID between video and audio modalities.
Omni-Full Attention. Previous works may employ separated spatial and temporal attention to reduce computational complexity, like UniForm [31]. However, as in CogVideoX [9], this separate attention mechanism requires extensive implicit information transmission, significantly increasing the learning complexity. Other works tailor two transformer towers for audio and video generation separately, e.g., AV-DiT, SyncFlow, JavisDiT, TAVGBench.However, they often adopt a multi-stage training approach, which is complex and resource-intensive. The two towers must first be pretrained separately, then finetuned together, increasing training time and resource consumption. To achieve more efficient training and more effective modality fusion, we employ the 3D text-video-audio hybrid full attention mechanism. As shown in Fig. 1, within the MM-DiT module, the hidden states of video, video-related text, audio-related text, and audio are first scaled and normalized, then concatenated together for the attention calculation.
$Q=Q_{V}\odot Q_{VT}\odot Q_{AT}\odot Q_{A},$ (3)
$K=K_{V}\odot K_{VT}\odot K_{AT}\odot K_{A},$ (4)
$V=V_{V}\odot V_{VT}\odot V_{AT}\odot V_{A},$ (5)
$Attn(Q,K,V)=Softmax(\frac{QK^{\top}}{\sqrt{d_{k}}})V,$ (6)
The attention values are then split into separate hidden states, which undergo scaling and normalization, residual connection, and feedforward, and subsequently fed to the next MM-DiT module. As a result, we achieve the unification of all input modalities within joint full-attention.
3.3 Multi-Task Progressive Training Strategy
Random Modality Masking. To learn generalizable and robust audio-visual representations for joint generation, we train the generative model with a broad spectrum of tasks. As a result, we propose to selectively adjust the mask of query and key for audio and video modalities. If we restrict the query and key to video embedding and video caption embedding, the model degenerates to a T2V model. Similarly, limiting the query and key to audio embedding and audio text embedding results in a T2A model. In this way, the model could not only handle joint generation, but also maintain the abilities of single-modality generation. Considering the scarcity of high-quality audio-video paired data, our method offers an alternative for training the T2VA model. We first pre-train Klear on T2V and T2A tasks, and then finetune our model on audio-video paired data to finally construct a T2VA model. The learning objectives for audio and video generation are in Eq. (7) and Eq. (8):
[…]
where $Mask_{a}$ is used to extract the audio token from the combined noise representation and $Mask_{v}$ is used to extract the vision tokens. In summary, $\mathcal{L}_{T2A}$ and $\mathcal{L}_{T2V}$ denote the single-modality tasks of T2A and T2V. To learn the generalizable and robust world knowledge of audio-visual correlation, we also incorporate several tasks of T2AV, I2V and I2AV. Consequently, the overall multi-task learning objective is as follows:
$\mathcal{L}_{overall}=\mathcal{L}_{T2A}+\mathcal{L}_{T2V}+\mathcal{L}_{T2AV}+\mathcal{L}_{I2V}+\mathcal{L}_{I2AV}$ (9)
Progressive Training Strategy. To efficiently train AV joint generation, we adopt a progressive multi-task learning framework with random modality masking applied throughout all stages:
Stage-I: Pre-training. We pretrain the model on the large-scale, multi-scene data corpus to acquire atomic generation capabilities across all tasks, including cross-modal semantic alignment, temporal synchronization, high-fidelity audio synthesis, and precise visual feature construction, which ensures basic abilities of both single modality generation and joint generation, and provides a solid foundation for subsequent post-training.
Stage-II: Specialized Post-training. We then specialize the model on its weaker abilities and tasks. Guided by evaluation metrics, we adaptively rebalance data distributions across scenarios and tasks to strengthen underperforming capabilities while preserving overall competency.
Stage-III: Quality-Refined Post-training. Finally, we fine-tune the model on the manually-curated, high-quality dataset to refine generation fidelity and enhance robustness in complex scenes, yielding improvements in perceptual realism and overall generation quality.
4 Dataset Construction
Figure 2: Overview of our Dataset Annotation Pipeline.
Dataset Overview. Our dataset comprises automatically annotated samples. The dataset contains single-speaker speech, multi-speaker speech, singing, and natural sound clips, with an overall post-filtering retention rate of 27%.
4.1 Dataset Filtering
Video Filtering and scene splitting. We first filter video quality by modeling dynamic quality (subject motion ratio, camera stability), static quality (sharpness, aesthetics, color saturation), content naturalness (no excessive effects/watermarks), and safety. We discard those videos with low resolution, low SNR/MOS, or over 20% silence. We then apply scene splitting to ensure each sample contains only one scene.
Audio Filtering and Post Processing. We filter audio data by removing samples with low SNR, MOS, abnormal clipping, distortion, or noise, ensuring less than 20% silence, high fidelity, and consistent formatting. We then assess audio–visual consistency, using Synchformer for temporal alignment and ImageBind for semantic alignment, ensuring high synchronization in both temporal and semantic dimensions.
4.2 Audio-Guided Data Splitting
We partition the dataset by audio type, separating vocal from non-vocal clips to form a sound split. From the vocal subset, we create singing, single-speaker speech, and multi-speaker speech splits, then apply dense captioning to each.
4.3 Dense Annotation and Integration
We annotate each split with specialized models for speech transcripts, audio captions, and video captions, including both meta information and detailed content. For speech and singing, we extract speaker attributes (e.g., gender, age), while the sound split receives only audio captions. We use Whisper-Large-v3, SenseVoice, and Qwen2.5-Omni for transcription, Qwen2.5-Omni and Gemini 2.5-Pro for audio captions, and a video expert model for detailed video labels. All annotations are merged into unified dense captions.
5 Experiments
In this section, we first present the experimental setup and implementation details (Sec. 5.1), then compare Klear with diverse baselines across multiple tasks (Sec. 5.2), complemented by qualitative results (Sec. 5.3). We further conduct ablations on the unified single-tower architecture, multi-task versus progressive training, and the role of 3D RoPE for native FPS (Sec. 5.4), collectively validating the effectiveness of our approach.
5.1 Experimental Setup
[…]
Baseline Methods. We select two canonical types of methods, including (1) Cascaded generation: these methods typically employ sequential T2V+V2A or T2A+A2V. Here, we employ AudioLDM2 [37] to perform the T2A task, while OpenSora [38] for T2V. (2) Joint generation. Recent works like JavisDiT [1], UniVerse-1 [2] and Ovi [3] employ a dual-tower architecture with dedicated interaction layers between them.
Evaluation Metrics. To evaluate the model’s generation capabilities, we test T2AV tasks on video, audio, and audio-video consistency. Following Universe-1, we use Verse-Bench for T2AV tasks. For video quality, we report the Motion Score (MS) based on RAFT [39] optical flow for dynamic realism and the Aesthetic Score (AS), a composite metric from MANIQA [40], aesthetic-predictor-v2-5 [41], and Musiq [42] for visual fidelity. Identity preservation is measured by ID Consistency (ID), calculated via DINOV3 [43] feature similarity. For audio quality, we use Fréchet Distance (FD) and KL Divergence on mel-spectrograms from PANNs [44]. Semantic alignment is evaluated using the CLAP score [45]. For synchronization, we report AV-A (Audio-Video Alignment) distance from Synchformer [46] and SyncNet Confidence (SNC) score [47] for lip sync. Global cross-modal alignment is measured by ImageBind (IB).
Table 1: Main comparisons of audio-visual joint generation.
Method Framework Video Audio TTS AV Consistency
MS $\uparrow$ AS $\uparrow$ ID $\uparrow$ FD $\downarrow$ KL $\downarrow$ CLAP $\uparrow$ WER $\downarrow$ AV-A $\downarrow$ SNC $\uparrow$ IB-Score $\uparrow$
AudioLDM+TemoTkn T2A+A2V 0.05 0.12 0.07 2.05 2.53 0.229 - 1.15 2.68 0.137
OpenSora+FoleyGen T2V+V2A 0.42 0.44 0.56 3.69 3.08 0.212 - 0.92 2.76 0.159
OpenSora+See&Hear 0.42 0.44 0.56 2.26 2.97 0.206 0.337 0.86 2.85 0.164
JavisDiT T2AV 0.18 0.36 0.22 1.95 5.17 0.228 0.256 0.92 3.94 0.231
SVG 0.40 0.41 0.25 1.55 3.62 0.080 - 0.72 4.07 0.206
Universe-1 0.20 0.47 0.25 1.55 1.25 0.160 0.180 0.98 3.92 0.198
Ovi 0.58 0.48 0.46 1.50 1.19 0.224 0.035 0.82 4.28 0.214
Klear (Ours) Unified T2AV 0.48 0.51 0.59 1.36 1.06 0.232 0.028 0.65 6.79 0.316
5.2 Comparison with Existing Methods
Klear demonstrates robust audio–video joint generation. In Table 1, Klear achieves state-of-the-art performance, surpassing the two prior methods by a large margin. Cascaded approaches perform poorly due to error accumulation and strong dependence on upstream generation quality, while existing joint models exhibit only moderate audio–video consistency and noticeable unimodal degradation. In contrast, Klear attains substantially better A/V consistency and synchronization, which we attribute to the unified single-tower architecture and omni full attention mechanism.
Klear effectively guarantees unimodal performance. We then assess unimodal performance. As shown in Table 1, Klear delivers high audio and video quality in joint generation—surpassing cascaded and joint baselines by 34% and 18%, respectively—while multitask training yields more generalizable representations. Moreover, Klear outperforms specialized T2A and T2V models on their respective tasks, indicating that leveraging complementary audio–visual knowledge strengthens unimodal representations and further improves single-modality generation quality.
Klear consistently maintains a performance advantage across multiple tasks. Table 1 provides a comprehensive evaluation of Klear across a broad suite of tasks—including TI2AV, TI2V, T2V, and T2A—with comparisons to task-specialized baselines. Klear delivers consistently strong results, matching or surpassing specialized state-of-the-art models.
Figure 3: Qualitative evaluation of audio-video joint generation across various aspects.
5.3 Qualitative Results
Lip-Sync Accuracy. Fig. 3 (a–b) highlights lip-sync performance. Klear achieves phoneme-level alignment between mouth movements and audio—covering mouth openings, lip–teeth shapes, and tongue positions, while Universe-1 and Ovi suffer from misalignment, delay, and clear audio–visual mismatch.
Emotional Expressiveness. From Fig. 3 (c), Klear generates characters with expressive emotions: facial cues (eyes, mouth curvature, muscle tension) are highly consistent with the audio’s affective tone (joy, sadness, excitement, lethargy), reflecting natural audio–visual fusion of prosody and dynamics. In contrast, Universe-1 and Ovi often produce distorted or emotional expressions.
Table 2: Comparison of different methods. The Dual Tower uses standard cross-attention, while the Single Tower utilizes our proposed Omni-Full Attention.
    Method Video Audio TTS Audio-Video Consistency
ID $\uparrow$ MOS $\uparrow$ CLAP $\uparrow$ WER $\downarrow$ DeSync $\downarrow$ Sync-conf $\uparrow$ IB $\uparrow$
    Dual Tower 0.62 62.02 0.139 0.675 1.163 3.762 0.126
    Single Tower 0.80 93.11 0.232 0.028 0.650 6.787 0.316
Singing and Rap Performance. As shown in Fig. 3, Klear yields performances where pitch, rhythm, and breath control are tightly aligned across audio and visuals—vibrato, melisma, and dynamic changes naturally match breathing patterns and facial expressions, consistent with human expectations of singing. In contrast, Universe-1 and Ovi show clear lip-sync failures, resulting in pronounced incongruity and reduced realism, especially for rap.
Audio–Visual Synchronization and Audio Overlapping. As shown in Fig. 3, Klear jointly generates background music and sound effects that are emotionally consistent with the video, with synchronized timing, realistic acoustics, and high fusion, thereby enhancing immersion. In contrast, baseline methods struggle to produce overlapping sounds and exhibit poor audio–visual consistency.
Image to Audio-Video. Fig. 3 presents TI2AV and TI2V results. Our model preserves high identity consistency with the input image while generating plausible camera motion and dynamics, whereas baselines exhibit identity drift, large visual discrepancies, and mechanical movements.
5.4 Ablation and Analysis
Architectural Effectiveness. To compare single- and dual-tower architectures for audio–video generation, we feed audio and video features into separate mm-DiT branches, with a randomly initialized audio tower due to the lack of a matching DiT backbone. Each block uses full attention and cross-attention for feature alignment. We also conduct an ablation with a pretrained audio tower. As shown in Table 2, the single-tower model outperforms the dual-tower variant in audio quality, video quality, and audio–video consistency, with visualizations confirming better cross-modal alignment. Although the pretrained audio tower converges faster, its performance is suboptimal due to a distribution mismatch with video features, hindering alignment.
Table 3: Ablation of multi-task masking, with arrows indicating optimization directions.
    Method Video Audio TTS Audio-Video Consistency
ID $\uparrow$ MOS $\uparrow$ CLAP $\uparrow$ WER $\downarrow$ DeSync $\downarrow$ Sync-conf $\uparrow$ IB $\uparrow$
    T2V 0.71 - - - - - -
    T2V+T2AV 0.76 88.181 0.188 0.044 0.895 5.024 0.201
    All Tasks(Ours) 0.80 93.106 0.232 0.028 0.650 6.787 0.316
Advantages of Multi-Task Masking. As shown in Table 3, for T2AV joint generation, our multi-task model significantly outperforms a counterpart trained solely on T2AV. This approach captures cross-modal audio–video correlations and complementary cues, outperforming video-only models on T2V and I2V. The unified multi-task training also produces robust, generalizable representations that scale well with data and compute, and the model generalizes effectively to I2AV and I2V, demonstrating high transferability.
Figure 4: Ablations of different training stages. Metrics include video, audio, TTS, and audio-video consistency, with arrows indicating optimization directions.
Gains from Progressive Training Strategy. As in Figure 4, the progressive training first equips the model with basic audio–video generation capabilities (a). In the post-training phase, the full model. The SP stage further boosts previously weak skills (e.g., IB score by 0.1), and post-training on high-quality data brings an additional overall gain over (b). Removing the entire progressive schedule instead causes a significant drop, confirming the effectiveness of our multi-stage training strategy.
6 Conclusion
We identify key failure modes in audio–video generation, such as asynchrony, lip–speech mismatch, and unimodal degradation, caused by suboptimal architectures, misaligned data, and single-task training. To address these, we propose Klear, a unified multi-task framework with a single-tower backbone and Omni-Full Attention for cross-modal interaction, alongside a progressive training strategy and an automated annotation pipeline. This pipeline produces a high-quality, annotated audio–video dataset for scalable training. Klear outperforms prior state-of-the-art on unimodal and joint audio-video benchmarks, becoming the superior model comparable to Veo 3. We hope this work could provide a clear direction and catalyze deeper research in audio–video generation.
7 Contributors
All contributors are listed as follows.
Jun Wang, Chunyu Qiang, Yuxin Guo, Yiran Wang, Feng Deng, Xijuan Zeng, Nan Li, Le Wang, Ming Wen, Zihan Li, Junjie Yan, Jingru Zhao, Zewen Song, Jiachen Zheng, Shiyao Wang, Xu Li, Zihao Ji, Zhongliang Liu, Xiaopeng Wang, Kang Yin, Teng Ma, Ziyu Zhang, Yuzhe Liang, Yushen Chen, Pengfei Cai, Haorui Zheng, Youjun Chen, Qianyue Hu, Jiahui Zhao, Yuejiao Wang, Liang Hou, Yuan Gao, Yan Zhou, Chen Zhang, Pengfei Wan.